# Latent Speech-Text Transformer

### v1

@@ 5 Results
Visualization of Word-Level Speech Patch Embeddings We use t-SNE (van der Maaten & Hinton, 2008) to project embeddings of representative word groups from the aligned-patching LST model (Fig. 4). Across different categories, embeddings of the same word consistently form tight clusters, while different words remain well separated. Each word forms its own cluster (e.g., he, she, they in pronouns; knife, scissors, sharpener in tools; boat, canoe, surfing in water-related terms). Related variants such as sail–sailing show stability under inflection, while semantically similar pairs like scissors–shears also appear nearby despite being distinct words. These qualitative patterns match quantitative results: within-word similarity is high (0.87), between-word similarity is much lower (0.43), and silhouette scores (0.65–0.68) (Rousseeuw, 1987) confirm well-separated clusters.
Scaling Trends. Table 5 summarizes results at both 1B and 7B scales, while Figure 5 provides the training curve of the 7B model on HellaSwag. Scaling consistently improves performance across all datasets. At 1B, LST already outperforms the baseline (e.g., 41.3 vs. 36.8 on SS, and 49.2 vs. 47.1 on TT). At 7B, the improvements persist: LST reaches 44.2/55.3 compared to the baseline’s 42.0/54.8. The figure further shows that LST exhibits a steeper growth curve over iterations, indicating more efficient utilization of larger capacity. Importantly, the 7B model remains far from convergence under the same processed token budget but with much fewer iterations, suggesting that extended training would likely amplify the advantage of LST and further widen the gap.
| Model | Batch | Iters | HellaSwag | |
| (M) | (k) | SS | TT | |
| Baseline (1B) | 0.5 | 200 | 36.8 | 47.1 |
| LST (1B) | 0.5 | 200 | 41.3 | 49.2 |
| Baseline (7B) | 4.0 | 25 | 42.0 | 54.8 |
| LST (7B) | 4.0 | 25 | 44.2 | 55.3 |
@@
Ablation on Patching Strategies Table 6 compares static and aligned patching. Aligned patching uses word boundaries from alignment, producing semantically coherent patches. We consider two variants: Align (sil sep.), keeping silence spans as separate patches, and Align (sil merged), merging them with adjacent words. Both outperform static patching at similar patch sizes—for instance, Align (sil sep.) reaches 60.3 on StoryCloze SS vs. 58.7 for static size 6, and Align (sil merged) scores 38.5 on HellaSwag SS vs. 37.2 for static size 9. Curriculum starts with Align (sil sep.) and gradually shifts to Static during training, retaining alignment benefits while matching the shorter-patch evaluation regime; it yields the strongest and most consistent results (e.g., 41.3 on HellaSwag SS). Overall, aligned patching better preserves semantics than static, and curriculum combines alignment supervision with static-style evaluation for the best performance. For completeness, we also report BPE-aligned patching experiments in Appendix A.5.
@@ 6 Related Work
LLMs using speech tokens. Early neural audio generation methods included direct auto-regressive generation of the speech waveform (van den Oord et al., 2016), or using adversarial approaches (Kong et al., 2020). Following this, textless NLP work (Lakhotia et al., 2021) showed that by using discrete speech tokens obtained from self-supervised speech encoders (CPC, wav2vec 2.0, HuBERT) as targets for language modeling, can enable fully spoken LLMs. AudioLM (Borsos et al., 2023) further uses hierarchical generation, first predicts semantic tokens, and subsequent stages predict fine-grained acoustic tokens from SoundStream (Zeghidour et al., 2021), to achieve both high audio quality as well as long-term consistency. In addition to augmenting semantic speech tokens with pitch and style tokens to explicitly model expressivity, SpiritLM (Nguyen et al., 2025) also introduced interleaving speech modeling with text-tokens. More recently, Moshi (Défossez et al., 2024) propose a hierarchical inner monologue method, that jointly predicts time-aligned text and acoustic tokens (with distilled semantic information), together with modeling multiple-stream audio for handling full-duplex audio dialogues. Finally, similar to scaling laws for text LLMs (Hoffmann et al., 2022), Cuervo & Marxer (2024) fit scaling law curves to predict the performance of spoken LLMs, and find that they scale upto three order of magnitude more slowly than text LLMs.
@@
| Model | Ave Patch Size | HellaSwag | StoryCloze | TopicStoryCloze | |||
@@
| LST (Static) | 4 | 40.5 | 48.8 | 58.2 | 69.4 | 86.2 | 95.1 |
| LST (Curriculum) | 5.8* 4 | 41.3 | 49.2 | 58.6 | 67.8 | 86.6 | 95.4 |
| LST (Align, sil sep.) | 5.8* | 39.9 | 49.3 | 60.3 | 69.9 | 85.7 | 95.3 |
| LST (Static) | 6 | 39.4 | 49.2 | 58.7 | 69.6 | 84.9 | 94.9 |
| LST (Static) | 9 | 37.2 | 49.4 | 57.5 | 69.7 | 84.7 | 95.9 |
| LST (Align, sil merged) | 9.4 | 38.5 | 49.0 | 58.8 | 69.7 | 86.9 | 96.0 |
@@
Speech Understanding Benchmarks. Going beyond measuring only acoustic and phonetic capabilities of speech models using scores such as ABX (Kahn et al., 2020), Nguyen et al. (2020) established the Zero Resource Speech Benchmark 2021, comprising datasets/metrics to evaluate lexical (sWUGGY), syntactic (sBLIMP) and lexical-semantic (sSIMI) capabilities of spoken LLMs. Since these benchmarks contrast between very short speech segments, we found that dynamic compute approaches such as ours, do not yield significant improvements. However, subsequently, Hassid et al. (2023) introduced the sStoryCloze and TopicStoryCloze datasets, which are story completion benchmarks in the speech modality measuring commonsense/understanding abilities of Spoken LLMs. We use these benchmarks in this paper, together with a speech version of the popular HellaSWAG textual benchmark, also measuring commonsense reasoning capabilities.
@@ 8 Conclusion
We presented the Latent Speech-Text Transformer (LST), a patching-based framework that compresses speech tokens into latent units for efficient and balanced multimodal training. Our experiments demonstrate that LST consistently outperforms baseline SpeechLLMs, with curriculum patching delivering the most robust gains across diverse datasets. By reducing speech–text imbalance, LST improves speech understanding while maintaining strong text performance, and its advantages grow further with model scaling. These findings highlight LST as a practical and scalable approach to bridging speech and text, offering improved efficiency, stronger cross-modal transfer, and greater robustness under varying training conditions.
@@ Appendix A Appendix
A.1 Interleaved Data Construction
@@
A.2 Hyperparameters
A.2.1 Optimization and Training Configuration
@@
A.2.2 Model Architecture
Table 7 summarizes the hierarchical architecture used in our experiments. The model consists of a shallow local encoder, a deep global transformer, and a moderately deep local decoder. The local modules operate with restricted attention windows to capture fine-grained context, while the global transformer uses block-causal attention with rotary position embeddings (RoPE) to handle long-range dependencies efficiently. This design balances local detail preservation with scalable long-context modeling.
| Module | Layers | Dim. | Heads | Notes |
| Local Encoder | 1 | 1024 | 16 | Local window = 512 |
| Global Transformer | 25 | 2048 | 16 | Block-causal; RoPE () |
| Local Decoder | 9 | 1024 | 16 | Local window = 512 |
A.3 Compute-Optimal Scaling
Figure 6 presents HellaSwag accuracy under compute-optimal training across model sizes ranging from 420M to 1.8B. For text, we follow the scaling rule of Hoffmann et al. (2022), allocating 20× the model size in training tokens. Speech uses half of the text tokens, preserving a 2:1 ratio of text to speech tokens. At the smallest scale (420M), LST already outperforms the baseline, reaching 29.2% vs. 28.4% for speech and 31.6% vs. 31.1% for text. These improvements compound with scale: at 1.8B, LST (Speech) attains 39.0% compared to 35.3%, while LST (Text) achieves 46.3% over 45.7%. Overall, LST provides consistent gains in both modalities, with advantages apparent from the earliest scale and amplified as model capacity increases.
We further analyze model preference using the NLL difference in Fig. 7, defined as the average gap between the negative log-likelihood of the correct option and that of the incorrect ones. More negative values indicate stronger separation, i.e., the model assigns lower NLL to the correct choice relative to distractors. LST consistently yields larger-magnitude gaps than the baseline across both text and speech. For speech (Fig. 7(a)), the gap expands from about to , while the baseline only improves from to . For text (Fig. 7(b)), the difference deepens from at 420M to at 1.8B, compared to the baseline’s to . Since NLL difference serves as a more stable indicator of model preference than accuracy alone, these results provide clearer evidence of the improvements achieved by LST as scale increases.
A.4 Effect of Speech Proportion
A.5 BPE-Aligned Patching
| Model | Ave Patch Size | HellaSwag | StoryCloze | TopicStoryCloze | |||
| (tokens) | SS | TT | SS | TT | SS | TT | |
| LST (Word Align) | 5.8* | 40.0 | 49.9 | 59.4 | 68.6 | 84.8 | 94.6 |
| LST (BPE Align) | 5.0* | 39.2 | 50.1 | 55.6 | 69.1 | 79.6 | 95.6 |
| LST (Word Curr.) | 5.84 | 41.5 | 49.5 | 57.9 | 68.9 | 86.8 | 95.1 |
| LST (BPE Curr.) | 5.04 | 41.3 | 48.6 | 59.1 | 67.1 | 86.5 | 95.4 |
A.6 Stability Analysis Across Tasks

### v2

@@ 5.3 Downstream Transfer and Analysis
We visualize word-level speech patch embeddings using t-SNE (van der Maaten and Hinton, 2008) (Fig. 5) from the aligned-patching LST model. Across different categories, embeddings of the same word consistently form tight clusters, while different words remain well separated. Each word forms its own cluster (e.g., he, she, they in pronouns; knife, scissors, sharpener in tools; boat, canoe, surfing in water-related terms). Related variants such as sail–sailing show stability under inflection, while semantically similar pairs like scissors–shears also appear nearby despite being distinct words. These qualitative patterns match quantitative results: within-word similarity is high (0.87), between-word similarity is much lower (0.43), and silhouette scores (0.65–0.68) (Rousseeuw, 1987) confirm well-separated clusters.
Ablation on Patching Strategies.
@@
Table 6 compares static and aligned patching. Aligned patching uses word boundaries from alignment, producing semantically coherent patches. We consider two variants: Align (sil sep.), keeping silence spans as separate patches, and Align (sil merged), merging them with adjacent words. Both outperform static patching at similar patch sizes—for instance, Align (sil sep.) reaches 60.3 on StoryCloze SS vs. 58.7 for static size 6, and Align (sil merged) scores 38.5 on HellaSwag SS vs. 37.2 for static size 9. Curriculum starts with Align (sil sep.) and gradually shifts to Static during training, retaining alignment benefits while matching the shorter-patch evaluation regime; it yields the strongest and most consistent results (e.g., 41.3 on HellaSwag SS). Overall, aligned patching better preserves semantics than static, and curriculum combines alignment supervision with static-style evaluation for the best performance. For completeness, we also report BPE-aligned patching experiments in Appendix A.7.
@@ 6 Related Work
LLMs using speech tokens. Early neural audio generation methods included direct auto-regressive generation of the speech waveform (van den Oord et al., 2016), or using adversarial approaches (Kong et al., 2020). Following this, textless NLP work (Lakhotia et al., 2021) showed that by using discrete speech tokens obtained from self-supervised speech encoders (CPC, wav2vec 2.0, HuBERT) as targets for language modeling, can enable fully spoken LLMs. AudioLM (Borsos et al., 2023) further uses hierarchical generation, first predicts semantic tokens, and subsequent stages predict fine-grained acoustic tokens from SoundStream (Zeghidour et al., 2021), to achieve both high audio quality as well as long-term consistency. In addition to augmenting semantic speech tokens with pitch and style tokens to explicitly model expressivity, SpiritLM (Nguyen et al., 2025) also introduced interleaving speech modeling with text-tokens. More recently, Moshi (Défossez et al., 2024) propose a hierarchical inner monologue method, that jointly predicts time-aligned text and acoustic tokens (with distilled semantic information), together with modeling multiple-stream audio for handling full-duplex audio dialogues. Finally, similar to scaling laws for text LLMs (Hoffmann et al., 2022), Cuervo and Marxer (2024) fit scaling law curves to predict the performance of spoken LLMs, and find that they scale upto three order of magnitude more slowly than text LLMs.
@@
| Patching Strategy | Ave Patch Size | HellaSwag | StoryCloze | TopicStoryCloze | |||
@@
| Static | 4 | 40.5 | 48.8 | 58.2 | 69.4 | 86.2 | 95.1 |
| Curriculum (sil sep.) | 5.8* 4 | 41.3 | 49.2 | 58.6 | 67.8 | 86.6 | 95.4 |
| Curriculum (sil merged) | 9.4 4 | 40.3 | 48.9 | 58.7 | 68.9 | 86.8 | 95.4 |
| Align (sil sep.) | 5.8* | 39.9 | 49.3 | 60.3 | 69.9 | 85.7 | 95.3 |
| Static | 6 | 39.4 | 49.2 | 58.7 | 69.6 | 84.9 | 94.9 |
| Static | 9 | 37.2 | 49.4 | 57.5 | 69.7 | 84.7 | 95.9 |
| Align, (sil merged) | 9.4 | 38.5 | 49.0 | 58.8 | 69.7 | 86.9 | 96.0 |
@@
Speech Understanding Benchmarks. Going beyond measuring only acoustic and phonetic capabilities of speech models using scores such as ABX (Kahn et al., 2020), Nguyen et al. (2020) established the Zero Resource Speech Benchmark 2021, comprising datasets/metrics to evaluate lexical (sWUGGY), syntactic (sBLIMP) and lexical-semantic (sSIMI) capabilities of spoken LLMs. Since these benchmarks contrast between very short speech segments, we found that dynamic compute approaches such as ours, do not yield significant improvements (see Appendix A.9). However, subsequently, Hassid et al. (2023) introduced the sStoryCloze and TopicStoryCloze datasets, which are story completion benchmarks in the speech modality measuring commonsense/understanding abilities of Spoken LLMs. We use these benchmarks in this paper, together with a speech version of the popular HellaSWAG textual benchmark, also measuring commonsense reasoning capabilities.
@@ 8 Conclusion
We presented the Latent Speech-Text Transformer (LST), a patch-based framework that aggregates speech tokens into latent units to improve computational efficiency and balance across modalities in multimodal language modeling. Across controlled-budget evaluations, scaling analyses, and downstream transfer experiments, LST consistently outperforms SpeechLLM baselines while reducing autoregressive sequence length and inference cost. Importantly, the gains arise from shortening the effective autoregressive sequence through latent speech patching, enabling more compute-efficient model scaling without sacrificing speech coverage or reconstruction quality. These findings highlight token-density imbalance as a key bottleneck in scaling spoken language models and suggest LST as a practical step toward compute-efficient unified speech–text foundation models.
@@ Appendix A Appendix
A.1 Data Construction and Licensing
A.1.1 Interleaved Data Construction
@@
A.1.2 Dataset Licensing
We summarize the licenses associated with the speech corpora used for pre-training.
Libri-Light. Derived from LibriVox audiobooks that are in the public domain.
The People’s Speech. Includes audio released under CC BY and CC BY-SA licenses.
Multilingual LibriSpeech (MLS). Distributed under CC BY 4.0.
Spotify Podcast Dataset. The dataset is described by Clifton et al. (Clifton et al., 2020b), available at https://arxiv.org/abs/2004.04270.
A.2 Additional Model Details
A.2.1 Local Patch Encoder and Decoder Details
We provide additional details on the local patch encoder and decoder that form the core representation transformation in LST.
Patch Encoder.
Following the BLT-style latent aggregation paradigm (Pagnoni et al., 2024), the patch encoder converts token-level speech representations into latent patch embeddings through alternating sliding-window self-attention and cross-attention layers. Within the encoder cross-attention, latent patch queries are initialized by pooling the hidden states of speech tokens within each local span. These pooled representations then act as queries, while the corresponding speech token features provide the keys and values. Self-attention operates over a restricted temporal window to preserve local coherence while controlling computational cost, after which cross-attention aggregates the token features into a smaller set of latent patch queries that summarize contiguous speech segments. Unlike BLT, LST: (i) applies patching only to speech-token spans determined by the strategy in Section 3.1, (ii) does not employ hash embeddings, as they showed no empirical benefit in our ablations. Additionally, LST does not rely on subword tokenization (e.g., BPE applied to speech tokens), which we find provides no performance benefit in our experiments.
Patch Decoder.
The patch decoder is implemented as a lightweight Transformer trained with next-token prediction (NTP) loss. Each decoder layer contains: (i) causal self-attention over a sliding window of the previously generated 512 tokens to ensure autoregressive consistency, and (ii) cross-attention where current token hidden states act as queries, while previously generated speech patches together with text tokens provide the keys and values. This allows token-level prediction to condition simultaneously on higher-level latent patch structure and textual context. This dual conditioning enables efficient generation while preserving long-range semantic and acoustic information encoded in the latent patches.
Intuition.
Conceptually, the local encoder–decoder pair performs a token latent patch token information bottleneck. The encoder compresses dense speech-token sequences into structured latent summaries, while the decoder re-expands them for autoregressive prediction. This mechanism is central to LST’s ability to reduce effective sequence length and computational cost without sacrificing linguistic or acoustic fidelity.
A.2.2 Model Architecture
Table 7 summarizes the hierarchical architecture used in our experiments. The model consists of a shallow patch encoder, a deep global transformer, and a moderately deep patch decoder. The local patch modules operate with restricted attention windows to capture fine-grained context, while the global transformer uses block-causal attention with rotary position embeddings (RoPE) to handle long-range dependencies efficiently. This design balances local detail preservation with scalable long-context modeling.
| Module | Layers | Dim. | Heads | Notes |
| Patch Encoder | 1 | 1024 | 16 | Patch window = 512 |
| Global Transformer | 25 | 2048 | 16 | Block-causal; RoPE () |
| Patch Decoder | 9 | 1024 | 16 | Patch window = 512 |
A.3 Optimization and Training Configuration
@@
A.4 Fixed-Token Scaling Across Model Sizes
Table 8 compares the baseline SpeechLLM and LST at 1B and 7B under an identical processed-token budget, ensuring fair cross-scale evaluation. At 1B, LST shows clear gains over the baseline (41.3 vs. 36.8 on SS and 49.2 vs. 47.1 on TT), indicating improved sample efficiency at smaller scale. At 7B, the advantage remains consistent, with LST achieving 44.2/55.3 compared to the baseline’s 42.0/54.8. These results demonstrate that the benefits of latent speech–text modeling persist across model scales and are not limited to low-capacity regimes.
| Model | Batch (M) | Iters (k) | HellaSwag | |
| SS | TT | |||
| Baseline (1B) | 0.5 | 200 | 36.8 | 47.1 |
| LST (1B) | 0.5 | 200 | 41.3 | 49.2 |
| Baseline (7B) | 4.0 | 25 | 42.0 | 54.8 |
| LST (7B) | 4.0 | 25 | 44.2 | 55.3 |
A.5 Stability Analysis Across Tasks
@@
A.6 Effect of Speech Proportion
A.7 BPE-Aligned Patching
| Model | Ave Patch Size | HellaSwag | StoryCloze | TopicStoryCloze | |||
| (tokens) | SS | TT | SS | TT | SS | TT | |
| LST (Word Align) | 5.8* | 40.0 | 49.9 | 59.4 | 68.6 | 84.8 | 94.6 |
| LST (BPE Align) | 5.0* | 39.2 | 50.1 | 55.6 | 69.1 | 79.6 | 95.6 |
| LST (Word Curr.) | 5.84 | 41.5 | 49.5 | 57.9 | 68.9 | 86.8 | 95.1 |
| LST (BPE Curr.) | 5.04 | 41.3 | 48.6 | 59.1 | 67.1 | 86.5 | 95.4 |
A.8 SC/TSC Evaluation Sets
We additionally evaluate on the publicly available speech versions of StoryCloze (SC) and TopicStoryCloze (TSC) (Hassid et al., 2023). The released TTS audio in these benchmarks contains noticeable synthesis artifacts and unnatural prosody, which may affect evaluation reliability. To provide more realistic acoustic conditions, we generate improved TTS renderings using the same text prompts and release them for reproducibility. Table 11 reports results on both the original and improved versions. Across all four settings, LST consistently outperforms the baseline, indicating that the observed gains are robust to TTS quality variations and do not depend on a specific synthesis condition.
| Model | SC (orig.) | SC (ours) | TSC (orig.) | TSC (ours) |
| Baseline | 58.0 | 59.1 | 78.4 | 87.5 |
| LST | 60.8 | 61.2 | 79.5 | 87.9 |
A.9 Fine-Grained Lexical and Syntactic Information
Although our main evaluation focuses on narrative understanding and commonsense reasoning, LST does not discard fine-grained lexical or syntactic information. The patch decoder preserves token-level detail through localized reconstruction, while aggregation operates only over short spans without collapsing phonetic distinctions. To verify this, we evaluate on sWUGGY and sBLIMP, which probe subword discrimination and syntactic sensitivity in spoken language modeling. As shown in Table 12, LST performs on par with the baseline, suggesting that latent patching preserves the linguistic cues required for low-level discrimination without introducing trade-offs against higher-level reasoning performance.
| Model | sWUGGY | sBLIMP |
| Baseline | 72.5 | 58.9 |
| LST | 72.8 | 59.0 |
A.10 Sequence Length, Compute, and Inference Scaling
Under the compute-controlled setting, sequence composition is determined by token-level statistics of interleaved speech–text data. The data follow a fixed word ratio of (speech : text), specified by the data construction. Because discrete speech units have higher temporal density than text tokens, this ratio corresponds in practice to approximately speech tokens and text tokens within the interleaved portion of the baseline sequence. LST applies speech patching, reducing speech tokens from to while leaving text tokens unchanged. The total sequence length therefore changes from to (). Given the quadratic dependence of transformer FLOPs on sequence length, this difference corresponds to an overall compute change of approximately .
During TTS inference, speech patching reduces the number of autoregressive acoustic generation steps by approximately . Since acoustic step count dominates generation cost, this leads to a substantial reduction in overall generation computation. During ASR inference, the reduced number of acoustic sequence elements decreases the effective encoder context length by a similar factor, yielding corresponding efficiency gains in acoustic encoding.
| Interleaved Speech Data (1:2) | Text Data | Total | ||
| Model | Speech units | Text units | Text units | Sequence length |
| Baseline | 0.77 | 0.23 | 2.00 | 3.00 |
| LST | 0.19 | 0.23 | 2.00 | 2.42 |
| FLOPs Reduction | – | |||
A.11 Aligner Quality Robustness
Curriculum patching relies on alignment boundaries primarily during the early training stage before transitioning to static patching. To assess robustness to alignment noise, we perturb Wav2Vec2+CTC boundaries by 20–40 ms and re-evaluate alignment-based patching. This perturbation causes at most 1% degradation in the setting most sensitive to boundary precision (pure alignment patching). Because curriculum patching depends less on precise boundaries and gradually removes alignment during training, these results indicate that its performance benefits are expected to persist even when using simpler or less accurate aligners.
A.12 LLM Usage.
We used large language models solely for polishing some sentences in this paper.